OpenStudy (anonymous):

The second hand on a clock is 2.5" long. How far does the tip of the second hand move in 3 minutes and 13 seconds?

OpenStudy (phi):

interesting question. First, if it went all the way round (1 full circle) how far did it go ?

OpenStudy (anonymous):

it did not go all the way around the clock.

OpenStudy (phi):

I know, but the first step could be figure out how big a circle it makes, if it did go all the way around

OpenStudy (phi):

Do you remember how to find the size of a circle, if you know its radius ?

OpenStudy (anonymous):

2*pi*r

OpenStudy (phi):

yes, with r = 2.5 inches (then length of the second hand) C (for circumference) = 2*pi*2.5= 5 pi (no need to replace pi with 3.14159.... yet) How long does it take the second hand to go all the way around ?

OpenStudy (anonymous):

43200?

OpenStudy (anonymous):

seconds

OpenStudy (phi):

43200 what ? Have you ever seen a second hand go around ? I think it takes exactly 60 seconds (or exactly 1 minute) to go around once.

OpenStudy (phi):

See the first paragraph http://en.wikipedia.org/wiki/Clock_face for a description of a clock with hands.

OpenStudy (anonymous):

is the answer 49.1?

OpenStudy (phi):

If the second hand goes around once in 1 minute, it traveled 5 pi inches Now to answer the question. How far does the tip of the second hand move in 3 minutes and 13 seconds? How far does it go in exactly 3 minutes ?

OpenStudy (phi):

If it goes 5*pi in 1 minute, how far does it go in 3 minutes ?

OpenStudy (anonymous):

15 pi

OpenStudy (phi):

ok, now for the 13 seconds 13 seconds is part of a minute. Can you change 13 seconds to a fraction of a minute ?

OpenStudy (anonymous):

yes i can. is the answer 1.7 inches??

OpenStudy (phi):

I would change 13 seconds to 13/60 of a minute or 0.217 minutes (roughly, as a decimal) the clock goes 5*pi*0.217 inches more or 1.085 pi more together with the 15 pi we get 15 pi + 1.085 pi = 16.085 pi (or we could have done 3 minutes and 13 seconds is 3.217 minutes, and total distance is 5*pi*3.217 which is 16.085*pi inches) now it's time to multiply by 3.14 (pi , roughly)

OpenStudy (anonymous):

Thank you so much for you help and persistence. This is a difficult subject for me. I wil be back later to work on this.

OpenStudy (phi):

Hopefully you get 16.085*3.14= 50.5 inches if you use the pi key on your calculator, you should find (3+13/60)*2*pi*2.5= 50.5272818